DogGirl27:

A rectangle’s length is three times its width, w. Its area is 243 square units. Which equation can be used to find the width of the rectangle? w2 = 3(243) 3w2 = 243 4w2 = 243 3w2 = 3(243)

GoodFellow:

quickly draw a figure of that rectangle & mark the said dimensions & try to write for area

GoodFellow:

drawing a figure, solves a question 50%,

Florisalreadytaken:

1) drawing a figure for it is just pointless. 2) where the literal Pear did alex get those numbers from?

GoodFellow:

@florisalreadytaken wrote:
1) drawing a figure for it is just pointless. 2) where the literal Pear did alex get those numbers from?
not pointless, it helps him at his level

GoodFellow:

at my level its just calculations by mind

Alex1513:

sorry mix up

Alex1513:

Width = w Length = 3w Area = 243 square units length *width = 243 3w * w = 243

Florisalreadytaken:

"A rectangle’s `length is three times its width`, w. Its `area is 243` square units. Which equation can be used to find the width of the rectangle?" having that info we can write: \[\color{lightskyblue}{w \times 3w}= 243 sq.^2\] \[ \color{lightskyblue}{3w^2} =243 sq.^2 \] which option is that?

DogGirl27:

@florisalreadytaken wrote:
"A rectangle’s `length is three times its width`, w. Its `area is 243` square units. Which equation can be used to find the width of the rectangle?" having that info we can write: \[\color{lightskyblue}{w \times 3w}= 243 sq.^2\] \[ \color{lightskyblue}{3w^2} =243 sq.^2 \] which option is that?
thank youuuu